# CLASSMix: Adaptive stain separation-based contrastive learning with pseudo labeling for histopathological image classification

### v1

@@ 4.6 Ablation studies
We conducted ablation studies to further validate our model’s H/E contrastive learning structure and identify the components contributing to the classification results.
One of the critical parts of our model is the H/E contrastive loss that provides a regularization term and helps to generate representative features. As shown in Table 2, removing the contrastive loss term from total loss led to a significant drop in the test accuracy. On the Utah ccRCC dataset, the test accuracy dropped from $94.92\%$ to $84.57\%$ in CLASS model. On TCGA ccRCC dataset, the test accuracy declined from $83.06\%$ to $74.46\%$. Choosing Hematoxylin and Eosin channels as two views also plays a crucial role. Instead of choosing Hematoxylin and Eosin as two channels, we did extra experiments by directly selecting two channels from Red, Green and Blue channels in RGB images to form two views in CLASS model. As a result, we saw obvious reduction in the test accuracies. In addition, splitting the original RGB images into Hematoxylin and Eosin channels is more logical and makes more sense in the aspect of pathology and the true nature of H&E images. The effect of adaptive stain separation was also verified. We used globally fixed stain separation matrix to replace adaptive stain separation. The CLASS model’s test accuracy dropped from $94.92\%$ to $92.43\%$ on Utah ccRCC dataset and dropped from $83.06\%$ to $79.89\%$ on TCGA ccRCC dataset. We further did ablation study on the augmentations on original RGB images. The augmentation proved beneficial and improved classification accuracy from $91.53\%$ to $92.43\%$ on dataset from Utah ccRCC dataset and from $75.67\%$ to $79.89\%$ on TCGA ccRCC dataset with fixed stain separation in CLASS model. However, the improvement was limited when adaptive stain separation was adopted. One possible reason is that the augmentation disturbed the color pattern and made adaptive stain separation produce less accurate Hematoxylin and Eosin results.
We finally analyzed the role of pseudo labeling using MixUp. Table 2 shows that, by comparing CLASSMix with CLASS, we saw a substantial improvement by adding MixUp augmentation, especially on the TCGA ccRCC dataset.
@@ 5 Conclusion
In this paper, we proposed a novel semi-supervised classification model called CLASSMix for histopathological image classification by adopting adaptive stain separation-based contrastive learning and pseudo labeling using MixUp. We provided newly annotated Utah ccRCC dataset and TCGA ccRCC dataset. Experiments on them have shown that our CLASSMix model consistently reached the best classification results compared to other state-of-the-art models. Our model demonstrates the capability to perform accurate patch-level classification at various resolutions with only rough annotations on approximately 30 WSIs in training. The code for our model is also publicly available.
In theory, the advantage of semi-supervised learning is the end-to-end training despite the sacrifice of conveniency compared to self-supervised learning. The self-supervised learning freezes the pre-trained encoders in final training with labeled data, which lacks the flexibility to fine-tune the whole model. However, unfreezing the encoders in final training leads to overfitting as there is no longer access to the large unlabeled dataset, especially when labeled data is limited.
@@ Appendix A. Detailed algorithm of Adaptive Stain Seperation
The letter C represents a specific channel, like Red, Green, or Blue channel. The value $I_{0,C}$ means the intensity of light before passing through the specimen, which is background intensity, and the value $I_{C}$ is the intensity of light after passing through the specimen. We can treat the $OD$ value as a measurement of the absorption of light on different channels. Zero values in all channels on $OD$ space means pure white background where there is no tissue absorbing light on that pixel. So for each RGB pixel, we use a $3\times 1$ vector $V_{OD}=(OD_{R},OD_{G},OD_{B})^{\top}$ to show the values on OD space.
Second, dimensionality reduction is applied. In an ideal situation of H&E staining, $V_{OD}$ is a linear combination of fixed Hematoxylin and Eosin unit vectors: $V_{OD}=\alpha_{H}V_{H}+\alpha_{E}V_{E}$, where $V_{H}$ and $V_{E}$ are $3\times 1$ unit vectors on 3D OD space. So all OD vectors should approximately fall into the same 2D plane for all pixels in one image. We use Principal Component Analysis (PCA) on 3D OD space to get a 2D plane formed by two covariance matrix’s eigenvectors with the largest two eigenvalues. Figure 3.a is an example of the distribution of OD vectors that map onto the 2D plane. The brightness shows the density of pixels. The x-axis is the direction of the eigenvector with the largest eigenvalue. The y-axis is the direction of the eigenvector with the second-largest eigenvalue. The third dimension would be treated as a residual part and be abandoned. The x and y axes are orthogonal because the eigenvectors of the covariance matrix are orthogonal to each other. By applying the PCA method, the 2D plane we found is more informative, increasing the possibility that the separated H and E images contain more information and details.
@@
The third step is to find stain vectors $V_{H}$ and $V_{E}$[37] assumes every pixel sample in OD space must exist between the two stain vectors, considering the fact that each component should be non-negative. However, noise can’t be fully avoided. As a result, we allow 1% of samples to be outside of $V_{H}$ and $V_{E}$ respectively. Moreover, pixel samples that are within a distance of 0.1 from original point are not taken into account as they contain less stain and are more easily influenced by noise. An example of $V_{H}$ and $V_{E}$ is shown in Figure 3.b. The positive directions of x-axis and y-axis were carefully chosen such that H unit vector falls into quadrant IV and E unit vector falls into quadrant I.
After acquiring unit vectors $V_{H}$ and $V_{E}$, we are able to reconstruct the transformation matrix. In the previous PCA process, we got three unit eigenvectors $V_{x}$, $V_{y}$, $V_{Residual}$, which are orthogonal to each other and ordered by their eigenvalues from maximum to minimum. The $V_{H}$ and $V_{E}$ on x-y 2D OD space can be written as $V_{H}=\cos\theta_{H}V_{x}+\sin\theta_{H}V_{y}$, $V_{E}=\cos\theta_{E}V_{x}+\sin\theta_{E}V_{y}$. The transformation matrix from H, E, Residual to RGB on OD space would be
@@
If we define $Mat_{RGB\rightarrow HERes\,on\,OD}$ as inverse matrix of $Mat_{HERes\rightarrow RGB\,on\,OD}$, then:
@@
The main advantages of this stain separation method are simplicity and efficiency. We are able to perform it without the need of model training or complex calculations. Moreover, finding optimum H vectors and E vectors helps to handle stain variations. The final normalization based on the calculated maximum value further mitigates the effect of variations on the brightness and hue of stains.
Table 3: Recall, precision and F-score of different classification models on Utah ccRCC dataset. Mean and standard deviation on test set are calculated based on 3 runs. Blue numbers show fully-supervised learning results. Brown numbers show transfer learning results. Green numbers show self-supervised learning results. Black numbers show semi-supervised learning results.
@@
ResNet(Frz) $0.8534\pm 0.0035$ $0.9326\pm 0.0035$ $0.7781\pm 0.0121$ $0.8633\pm 0.0056$ $0.8568\pm 0.0018$
ViT(Frz) $0.8486\pm 0.0089$ $0.9415\pm 0.0073$ $0.7241\pm 0.0193$ $0.7861\pm 0.0088$ $0.8251\pm 0.0015$
BT $0.8987\pm 0.0466$ $0.9810\pm 0.0075$ $0.8612\pm 0.0268$ $0.8849\pm 0.0708$ $0.9065\pm 0.0121$
SwAV $0.9363\pm 0.0205$ $0.9537\pm 0.0126$ $0.8946\pm 0.0380$ $0.8640\pm 0.0542$ $0.9122\pm 0.0067$
MoCoV2 $0.9586\pm 0.0081$ $0.9426\pm 0.0042$ $0.7352\pm 0.0068$ $0.8575\pm 0.0204$ $0.8735\pm 0.0021$
ViT-DINO $0.9734\pm 0.0089$ $0.9950\pm 0.0029$ $0.9632\pm 0.0097$ $0.7529\pm 0.0411$ $0.9211\pm 0.0086$
@@
ResNet(Frz) $0.9902\pm 0.0004$ $0.4538\pm 0.0090$ $0.3983\pm 0.0088$ $0.8628\pm 0.0091$ $0.6762\pm 0.0006$
ViT(Frz) $0.9790\pm 0.0012$ $0.5345\pm 0.0078$ $0.3128\pm 0.0158$ $0.7876\pm 0.0077$ $0.6535\pm 0.0048$
BT $0.9868\pm 0.0075$ $0.8436\pm 0.0464$ $0.5036\pm 0.0831$ $0.7130\pm 0.1271$ $0.7618\pm 0.0442$
SwAV $0.9875\pm 0.0030$ $0.8012\pm 0.0868$ $0.6008\pm 0.0612$ $0.8228\pm 0.0980$ $0.8031\pm 0.0306$
MoCoV2 $0.9829\pm 0.0044$ $0.7937\pm 0.0255$ $0.5673\pm 0.0307$ $0.9063\pm 0.0246$ $0.8126\pm 0.0186$
ViT-DINO $0.9729\pm 0.0048$ $0.9479\pm 0.0228$ $0.7767\pm 0.0840$ $0.8782\pm 0.0295$ $0.8939\pm 0.0294$
@@
ResNet(Frz) $0.9167\pm 0.0019$ $0.6104\pm 0.0074$ $0.5268\pm 0.0062$ $0.8630\pm 0.0023$ $0.7292\pm 0.0017$
ViT(Frz) $0.9091\pm 0.0048$ $0.6819\pm 0.0080$ $0.4365\pm 0.0118$ $0.7868\pm 0.0014$ $0.7036\pm 0.0059$
BT $0.9401\pm 0.0222$ $0.9066\pm 0.0235$ $0.6318\pm 0.0565$ $0.7811\pm 0.0512$ $0.8149\pm 0.0290$
SwAV $0.9611\pm 0.0099$ $0.8687\pm 0.0476$ $0.7164\pm 0.0297$ $0.8383\pm 0.0264$ $0.8462\pm 0.0196$
MoCoV2 $0.9706\pm 0.0020$ $0.8616\pm 0.0151$ $0.6402\pm 0.0220$ $0.8809\pm 0.0044$ $0.8383\pm 0.0097$
ViT-DINO $0.9731\pm 0.0023$ $0.9707\pm 0.0107$ $0.8586\pm 0.0544$ $0.8097\pm 0.0117$ $0.9030\pm 0.0177$
@@
CLASSMix $0.9602\pm 0.0106$ $0.9416\pm 0.0037$ $0.8407\pm 0.0050$ $0.7864\pm 0.0479$ $0.8822\pm 0.0150$
Table 4: Recall, precision and F-score of different classification models on TCGA ccRCC dataset. Mean and standard deviation on test set are calculated based on 3 runs. Blue numbers show fully-supervised learning results. Brown numbers show transfer learning results. Green numbers show self-supervised learning results. Black numbers show semi-supervised learning results.
@@
ResNet(Frz) $0.7239\pm 0.0054$ $0.8837\pm 0.0031$ $0.5876\pm 0.0141$ $0.7317\pm 0.0042$
ViT(Frz) $0.7383\pm 0.0072$ $0.9233\pm 0.0038$ $0.5333\pm 0.0212$ $0.7316\pm 0.0046$
BT $0.7906\pm 0.0274$ $0.9650\pm 0.0029$ $0.6688\pm 0.0681$ $0.8081\pm 0.0179$
SwAV $0.7462\pm 0.0577$ $0.9436\pm 0.0227$ $0.6271\pm 0.0695$ $0.7723\pm 0.0361$
MoCoV2 $0.8377\pm 0.0545$ $0.9633\pm 0.0093$ $0.6152\pm 0.0072$ $0.8054\pm 0.0175$
ViT-DINO $0.7969\pm 0.0962$ $0.9782\pm 0.0049$ $0.5531\pm 0.0179$ $0.7760\pm 0.0364$
@@
ResNet(Frz) $0.5883\pm 0.0071$ $0.9391\pm 0.0016$ $0.5735\pm 0.0188$ $0.7003\pm 0.0044$
ViT(Frz) $0.6919\pm 0.0112$ $0.9093\pm 0.0011$ $0.7886\pm 0.0154$ $0.7966\pm 0.0079$
BT $0.8824\pm 0.0415$ $0.9324\pm 0.0072$ $0.7535\pm 0.0809$ $0.8561\pm 0.0163$
SwAV $0.8290\pm 0.0886$ $0.9401\pm 0.0106$ $0.5345\pm 0.0915$ $0.7679\pm 0.0394$
MoCoV2 $0.8391\pm 0.0288$ $0.9362\pm 0.0140$ $0.8759\pm 0.0072$ $0.8837\pm 0.0061$
ViT-DINO $0.8722\pm 0.0096$ $0.9214\pm 0.0199$ $0.9432\pm 0.0085$ $0.9123\pm 0.0070$
@@
ResNet(Frz) $0.6490\pm 0.0029$ $0.9106\pm 0.0023$ $0.5803\pm 0.0132$ $0.7133\pm 0.0039$
ViT(Frz) $0.7143\pm 0.0093$ $0.9162\pm 0.0016$ $0.6361\pm 0.0151$ $0.7554\pm 0.0040$
BT $0.8331\pm 0.0058$ $0.9484\pm 0.0051$ $0.7044\pm 0.0267$ $0.8286\pm 0.0117$
SwAV $0.7815\pm 0.0275$ $0.9417\pm 0.0062$ $0.5768\pm 0.0825$ $0.7667\pm 0.0362$
MoCoV2 $0.8370\pm 0.0128$ $0.9494\pm 0.0029$ $0.7227\pm 0.0060$ $0.8364\pm 0.0070$
ViT-DINO $0.8310\pm 0.0563$ $0.9488\pm 0.0086$ $0.6971\pm 0.0130$ $0.8257\pm 0.0258$

### v4

@@ 4.6 Ablation studies
We conducted ablation studies to further analyze our model and identify the components contributing to the classification results. One of the critical parts of our model is the H/E contrastive loss which provides a regularization term and helps to generate representative features. As shown in Table 2, removing the contrastive loss term led to a significant drop in the test accuracy. On the Utah ccRCC dataset, the test accuracy dropped from $95.35\%$ to $90.92\%$ for the CLASS-M model and from $94.92\%$ to $84.57\%$ for the CLASS model. On TCGA ccRCC dataset, the test accuracy declined from $92.13\%$ to $89.70\%$ for the CLASS-M model and from $83.06\%$ to $74.46\%$ for the CLASS model.
Choosing Hematoxylin and Eosin channels as two views also plays a crucial role. Instead of choosing Hematoxylin and Eosin as two channels, we experimented with directly selecting two channels from Red, Green and Blue channels in RGB images to form two views in the CLASS model. As a result, we saw large reductions in test accuracies. This is expected due to the independence requirement in co-training which Hematoxylin and Eosin fulfill to a large extent, but Red/Green/Blue channels do not. In addition, splitting the original RGB images into Hematoxylin and Eosin channels is motivated from the pathology point of view and the true nature of H&E images.
The effect of adaptive stain separation was also verified. We used globally fixed stain separation matrix to replace adaptive stain separation. The CLASS-M model’s test accuracy dropped from $95.35\%$ to $93.97\%$ on Utah ccRCC dataset and dropped from $92.13\%$ to $90.97\%$ on TCGA ccRCC dataset. We further did ablation study on the augmentations on original RGB images. The augmentations proved beneficial and improved classification accuracy from $94.41\%$ to $95.35\%$ on Utah ccRCC dataset and from $91.21\%$ to $92.13\%$ on TCGA ccRCC dataset for the CLASS-M model. We finally analyzed the role of pseudo-labeling using MixUp. Table 2 shows that, by comparing CLASS-M with CLASS, we saw a substantial improvement by adding MixUp augmentation, especially on the TCGA ccRCC dataset.
@@ 5 Conclusion
In this paper, we proposed a novel semi-supervised model named CLASS-M for histopathological image classification by introducing adaptive stain separation-based contrastive learning and adopting pseudo-labeling using MixUp. We provided the newly annotated Utah ccRCC dataset and TCGA ccRCC dataset. Experiments have shown that our CLASS-M model consistently reached the best classification results compared to other state-of-the-art models. Our model demonstrates the capability to perform accurate patch-level classification at various resolutions with only rough annotations on approximately 30 WSIs in training. We demonstrated that CLASS-M outperforms state-of-the-art general semi-supervised computer vision models. The code for our model is also publicly available.
The advantage of semi-supervised learning is the end-to-end training despite the sacrifice of conveniency compared to self-supervised learning. Self-supervised learning freezes the pre-trained encoders in final training with labeled data, which lacks the flexibility to fine-tune the whole model. However, unfreezing the encoders in final training leads to overfitting as there is no longer access to the large unlabeled dataset, especially when labeled data is limited. We demonstrated that our semi-supervised model is able to outperform self-supervised learning models.
@@
CRediT author statement
Bodong Zhang: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Visualization. Hamid Manoochehri: Software, Formal analysis, Visualization. Man Minh Ho: Software, Validation, Investigation. Fahimeh Fooladgar: Software, Validation, Investigation. Yosep Chong: Writing - Review & Editing. Beatrice S. Knudsen: Conceptualization, Resources, Data Curation, Writing - Review & Editing, Supervision, Project administration, Funding acquisition. Deepika Sirohi: Resources, Data Curation. Tolga Tasdizen: Conceptualization, Resources, Writing - Review & Editing, Supervision, Project administration, Funding acquisition.
Availability of research data/code
Our related research data and code is publicly available at github.com/BzhangURU/Paper_CLASS-M/tree/main
@@ Appendix A. Detailed algorithm of Adaptive Stain Seperation
The letter C represents a specific channel such as Red, Green, or Blue channel. The value $I_{0,C}$ refers to the intensity of light before passing through the specimen, which is background intensity, and the value $I_{C}$ is the intensity of light after passing through the specimen. We can treat the $OD$ value as a measurement of the absorption of light on different channels. Zero values in all channels on $OD$ space correspond to cases of pure white background where there is no tissue absorbing light on that pixel. For each RGB pixel, we use a $3\times 1$ vector $V_{OD}=(OD_{R},OD_{G},OD_{B})^{\top}$ to show the values on OD space.
Second, dimensionality reduction is applied. In an ideal situation of H&E staining, $V_{OD}$ is a linear combination of fixed Hematoxylin and Eosin unit vectors: $V_{OD}=\alpha_{H}V_{H}+\alpha_{E}V_{E}$, where $V_{H}$ and $V_{E}$ are $3\times 1$ unit vectors on 3D OD space. Therefore, all OD vectors should approximately fall into the same 2D plane for all pixels in one image. We use Principal Component Analysis (PCA) on 3D OD space to get a 2D plane formed by two covariance matrix’s eigenvectors with the largest two eigenvalues. Figure 4.a is an example of the distribution of OD vectors that map onto the 2D plane. The brightness shows the density of pixels. The x-axis is the direction of the eigenvector with the largest eigenvalue. The y-axis is the direction of the eigenvector with the second-largest eigenvalue. The third dimension can be treated as a residual part and abandoned. The x and y axes are orthogonal because the eigenvectors of the covariance matrix are orthogonal to each other. By applying the PCA method, the 2D plane we find is more informative, ensuring that the separated H and E images contain critical information and details.
@@
The third step is to find stain vectors $V_{H}$ and $V_{E}$.  [40] assumes every pixel sample in OD space must exist between the two stain vectors, considering the fact that each component should be non-negative. However, noise can’t be fully avoided. As a result, we allow 1% of samples to be outside of $V_{H}$ and $V_{E}$ respectively. Moreover, pixel samples that are within a distance of 0.1 from origin are not taken into account as they contain less stain and are more easily influenced by noise. An example of $V_{H}$ and $V_{E}$ is shown in Figure 4.b. The positive directions of x-axis and y-axis were carefully chosen such that H unit vector falls into quadrant IV and E unit vector falls into quadrant I.
After acquiring unit vectors $V_{H}$ and $V_{E}$, we are able to reconstruct the transformation matrix. In the previous PCA process, we computed three unit eigenvectors $V_{x}$, $V_{y}$, $V_{Residual}$, which are orthogonal to each other and ordered by their eigenvalues from maximum to minimum. The $V_{H}$ and $V_{E}$ on x-y 2D OD space can be written as $V_{H}=\cos\theta_{H}V_{x}+\sin\theta_{H}V_{y}$, $V_{E}=\cos\theta_{E}V_{x}+\sin\theta_{E}V_{y}$. The transformation matrix from H, E, Residual to RGB on OD space is found as
@@
If we define $Mat_{RGB\rightarrow HERes\,on\,OD}$ as the inverse of $Mat_{HERes\rightarrow RGB\,on\,OD}$, then:
@@
The main advantages of this stain separation method are simplicity and efficiency. We are able to perform it without the need of model training or complex calculations. Moreover, finding optimum H vectors and E vectors helps to handle stain variations. The final normalization based on the calculated maximum value further mitigates the effect of variations on the brightness and strength of stains.
Table 3: Recall, precision and F-score of different classification models on Utah ccRCC dataset. Mean and standard deviation on test set are calculated based on 3 runs. Blue numbers show fully-supervised learning results. Green numbers show self-supervised learning results. Black numbers show semi-supervised learning results.
@@
BT $0.9685\pm 0.0030$ $0.9693\pm 0.0026$ $0.9169\pm 0.0083$ $0.8820\pm 0.0097$ $0.9342\pm 0.0043$
SwAV $0.9413\pm 0.0049$ $0.9889\pm 0.0010$ $0.9674\pm 0.0039$ $0.8571\pm 0.0242$ $0.9387\pm 0.0059$
MoCo v3 $0.9631\pm 0.0068$ $0.9755\pm 0.0010$ $0.9306\pm 0.0089$ $0.8874\pm 0.0082$ $0.9391\pm 0.0060$
ViT-DINO $0.9017\pm 0.0230$ $0.9933\pm 0.0033$ $0.8946\pm 0.0567$ $0.8315\pm 0.0357$ $0.9053\pm 0.0113$
@@
BT $0.9924\pm 0.0008$ $0.9021\pm 0.0100$ $0.7107\pm 0.0134$ $0.8666\pm 0.0277$ $0.8679\pm 0.0076$
SwAV $0.9855\pm 0.0031$ $0.9589\pm 0.0018$ $0.6785\pm 0.0070$ $0.7442\pm 0.0212$ $0.8418\pm 0.0029$
MoCo v3 $0.9860\pm 0.0010$ $0.9149\pm 0.0132$ $0.8037\pm 0.0276$ $0.8271\pm 0.0279$ $0.8829\pm 0.0171$
ViT-DINO $0.9803\pm 0.0023$ $0.7384\pm 0.0237$ $0.6783\pm 0.1003$ $0.6655\pm 0.0961$ $0.7657\pm 0.0358$
@@
BT $0.9803\pm 0.0018$ $0.9345\pm 0.0044$ $0.8007\pm 0.0079$ $0.8741\pm 0.0167$ $0.8974\pm 0.0060$
SwAV $0.9628\pm 0.0011$ $0.9737\pm 0.0014$ $0.7976\pm 0.0062$ $0.7963\pm 0.0020$ $0.8826\pm 0.0012$
MoCo v3 $0.9744\pm 0.0039$ $0.9442\pm 0.0074$ $0.8624\pm 0.0197$ $0.8561\pm 0.0180$ $0.9093\pm 0.0122$
ViT-DINO $0.9393\pm 0.0132$ $0.8469\pm 0.0146$ $0.7662\pm 0.0428$ $0.7381\pm 0.0721$ $0.8226\pm 0.0235$
@@
CLASS-M $0.9602\pm 0.0106$ $0.9416\pm 0.0037$ $0.8407\pm 0.0050$ $0.7864\pm 0.0479$ $0.8822\pm 0.0150$
Table 4: Recall, precision and F-score of different classification models on TCGA ccRCC dataset. Mean and standard deviation on test set are calculated based on 3 runs. Blue numbers show fully-supervised learning results. Green numbers show self-supervised learning results. Black numbers show semi-supervised learning results.
@@
BT $0.8390\pm 0.0322$ $0.9003\pm 0.0145$ $0.5833\pm 0.1345$ $0.7742\pm 0.0492$
SwAV $0.9542\pm 0.0029$ $0.9714\pm 0.0011$ $0.5393\pm 0.0029$ $0.8217\pm 0.0005$
MoCo v3 $0.8479\pm 0.0329$ $0.9559\pm 0.0078$ $0.5609\pm 0.0030$ $0.7882\pm 0.0093$
ViT-DINO $0.9222\pm 0.0511$ $0.9857\pm 0.0047$ $0.4850\pm 0.0577$ $0.7976\pm 0.0215$
@@
BT $0.7161\pm 0.0457$ $0.9398\pm 0.0256$ $0.6184\pm 0.1002$ $0.7581\pm 0.0162$
SwAV $0.8597\pm 0.0039$ $0.9581\pm 0.0011$ $0.9211\pm 0.0055$ $0.9130\pm 0.0025$
MoCo v3 $0.7902\pm 0.0241$ $0.9404\pm 0.0100$ $0.8681\pm 0.0102$ $0.8662\pm 0.0082$
ViT-DINO $0.9420\pm 0.0244$ $0.9374\pm 0.0100$ $0.9210\pm 0.0125$ $0.9335\pm 0.0068$
@@
BT $0.7718\pm 0.0266$ $0.9194\pm 0.0098$ $0.5849\pm 0.0325$ $0.7587\pm 0.0195$
SwAV $0.9045\pm 0.0009$ $0.9647\pm 0.0006$ $0.6803\pm 0.0032$ $0.8498\pm 0.0014$
MoCo v3 $0.8174\pm 0.0050$ $0.9481\pm 0.0012$ $0.6814\pm 0.0021$ $0.8156\pm 0.0022$
ViT-DINO $0.9314\pm 0.0264$ $0.9609\pm 0.0058$ $0.6337\pm 0.0463$ $0.8420\pm 0.0200$
@@
Appendix C. Visualization of our CLASS-M model predictions.
Please check Figure 5 and Figure 6 for visualization of our CLASS-M model’s patch-level predictions on WSIs from test set for both datasets.
Figure 5: Visualization of our CLASS-M model’s predictions on Utah ccRCC WSIs from test set. The polygons show the annotations on WSIs. (Green: Normal/Benign, Yellow: Low risk cancer, Red: High risk cancer, Dark Grey: Necrosis) Each polygon labeled as low/high risk cancer marks the region of a certain cancer growth pattern. The heatmaps show the model’s predictions on all foreground tiles when the predictions are not Normal/Benign. The strength of color shows the confidence of prediction.
Figure 6: Visualization of our CLASS-M model’s predictions on TCGA ccRCC WSIs from test set. The polygons show the annotations on WSIs. (Green: Normal/Benign, Orange: Cancer, Dark Grey: Necrosis) Each polygon labeled as cancer marks the region of a certain cancer growth pattern. The heatmaps show the model’s predictions on all foreground tiles when the predictions are not Normal/Benign. The strength of color shows the confidence of prediction.